OpenStudy (anonymous):

the average of y, 2y, and 3y is 12, then y=

OpenStudy (anonymous):

To get the average, you sum the three terms and divide by three. Equate your result with 12 and solve for y.

OpenStudy (aayushi.somani):

yeah y= 6 ..!

OpenStudy (anonymous):

(y + 2y + 3y) / 3 = (6y) / 3 = 2y This is 12. Divide by two to get the result y = 6.

OpenStudy (timmyk14):

Ok, average is another name for the mean of the data set... So y + 2y + 3y = 12 When you find the average (mean) you, in the end, divide by the number of numbers in the data set...3 12 * 3 = 36 So, y + 2y + 3y actually equals 36 Now we solve the equation... 6y = 36 (after combining like terms) --- --- 6. 6 ----------- Y = 6 Hope I helped (^_^)

OpenStudy (anonymous):

What is average? Sum of your numbers devided by how many numbers there are.